### v1

# CFSP: An Efficient Structured Pruning Framework for LLMs with Coarse-to-Fine Activation Information

Abstract
The colossal parameters and computational overhead of Large Language Models (LLMs) challenge their real-world applications. Network pruning, which targets unstructured or structured sparsity by removing redundant parameters, has recently been explored for LLM acceleration. Existing LLM pruning works focus on unstructured pruning, which typically requires special hardware support for a practical speed-up. In contrast, structured pruning can reduce latency on general devices. However, it remains a challenge to perform structured pruning efficiently and maintain performance, especially at high sparsity ratios. To this end, we introduce an efficient structured pruning framework named CFSP, which leverages both Coarse (interblock) and Fine-grained (intrablock) activation information as an importance criterion to guide pruning. The pruning is highly efficient, as it only requires one forward pass to compute feature activations. Specifically, we first allocate the sparsity budget across blocks based on their importance and then retain important weights within each block. In addition, we introduce a recovery fine-tuning strategy that adaptively allocates training overhead based on coarse-grained importance to further improve performance. Experimental results demonstrate that CFSP outperforms existing methods on diverse models across various sparsity budgets. Our code will be available at https://github.com/wyxscir/CFSP.
CFSP: An Efficient Structured Pruning Framework for LLMs with Coarse-to-Fine Activation Information
Yuxin Wang [^1] footnotemark: 1, Minghua Ma [^1] footnotemark: 1, Zekun Wang1thanks: Equal Contribution.thanks: Corresponding Author., Jingchang Chen1, Huiming Fan1 Liping Shan2, Qing Yang2, Dongliang Xu2, Ming Liu122footnotemark: 2, Bing Qin1 1Harbin Institute of Technology, Harbin, China 2Du Xiaoman Science Technology Co., Ltd, Beijing, China {yuxinwang,mhma,zkwang,jcchen,hmfan,mliu,qinb}@ir.hit.edu.cn
1 Introduction
Although scaling up Large Language Models (LLMs) brings remarkable performance (Brown et al., 2020; OpenAI, 2023; Gemini Team et al., 2023; Meta, 2024; DeepSeek-AI et al., 2024; Yang et al., 2024a), increasing parameters brings more computations and memory consumption, posing a significant challenge of deploying in practical applications. To address this, various model compression methods for LLMs are proposed (Dettmers et al., 2022; Frantar et al., 2022; Lin et al., 2023; Muralidharan et al., 2024). Existing LLM pruning work (Frantar and Alistarh, 2023; Sun et al., 2023; Xu et al., 2024a; Zhang et al., 2024b) focuses mainly on unstructured or semi-structured sparsity. However, these paradigms require specific hardware to achieve practical acceleration.
In contrast, structured pruning, which imposes structured sparsity by removing groups of consecutive parameters (Louizos et al., 2017; Wang et al., 2020; Xia et al., 2022), is more hardware-friendly on general devices. However, there are some challenges involved in existing structured pruning methods for LLMs: (1) They typically introduce learnable masks to search (Xia et al., 2023; Dery et al., 2024) or utilize gradients to guide pruning (Ma et al., 2023; Zhang et al., 2023a). Unfortunately, they require significant computational overhead, especially for large-scale (e.g., 70B) models. (2) It is also worth noting that they usually assign a uniform sparsity budget per block, which is suboptimal since LLM blocks have different significance in the representation functionality (Gromov et al., 2024a). Moreover, they usually involve a recovery fine-tuning with Low-Rank Adapter (LoRA) (Hu et al., 2022) to enhance pruned models, which also introduce training overhead and overlook the varying importance of blocks.
Figure 1: Illustration of our proposed CFSP framework. (a) Pruning with coarse (interblock) to fine (intrablock) activation information guidance. (b) Recovery fine-tuning with importance-guided allocation, where the rank sizes of each component are determined by coarse-grained importance.
To this end, we propose CFSP (shown in Figure 1), an efficient structural pruning framework for LLM that takes advantage of coarse to fine-grained activation information to guide pruning. Specifically, we employ activation as the importance criterion, which is calculated for blocks (coarse-grained) and the weights within each block (fine-grained) in a single forward pass. For each block, we measure its saliency of transformations based on the angular distance of the input and output representations. Then, we utilize this criterion as coarse-grained importance to assign the sparsity budget. Finally, for weights within each block, we take the product of their relative activations and weights as a fine-grained criterion to remove redundant parts. Since existing work typically performs recovery fine-tuning with LoRA to further improve performance, we propose a more efficient recovery method that leverages coarse-grained importance to adaptively allocate additional trainable parameters: the pruned models can achieve comparable performance while utilizing less recovery data.
Our contributions can be summarized as:
  • We propose an efficient coarse-to-fine importance criterion for identifying redundant structures for pruning, which takes only a few minutes [^1] Details of time cost are shown in Table 7 in Appendix. to complete on various models.
  • We introduce an efficient recovery fine-tuning method that adaptively assigns additional trainable parameters based on coarse-grained importance score.
  • Extensive experimental results indicate that CFSP surpasses existing methods across various models at different sparsity levels, demonstrating promising performance on challenging tasks even at high sparsity levels.
2 Methodology
The overview of CFSP framework is shown in Figure 1. We first introduce our preliminary analysis in Section 2.1, then give details of our pruning criterion and procedure in Section 2.2. Finally, we introduce the proposed importance-guidance recovery strategy in Section 2.3.
2.1 Preliminaries
The Transformer block (Vaswani et al., 2017) consists of multi-head attention (MHA) and feedforward network (FFN). We analyze the computational overhead and the sparsity of them. As shown in Figure 2, the parameter size and MAC of FFN are significantly larger than those of MHA. In addition, we observe that pruning MHA leads to a significant performance drop with only 10% sparsity, while pruning FFN has a more stable performance even with 50% sparsity, showing that the FFN module has a higher structural sparsity (Zhang et al., 2022) and is more friendly to structured pruning (Gunter et al., 2024). Thus, in this work, we focus on pruning the intermediate dimension of FFN.
2.2 CFSP Framework
CFSP takes activations as an importance criterion to identify redundant parts of LLMs for the following reasons: (1) Activations can be obtained with a single forward pass, resulting in significantly lower overhead compared to other metrics. (2) As pointed out in previous studies (Sun et al., 2023; Lin et al., 2023), parameter weights corresponding to larger activation magnitudes are more salient since they process more important features.
The feature activations are calculated on a small number (e.g., 128) of calibration samples. We further incorporate coarse- and fine-grained importance for sparsity allocation and weight pruning.
Figure 2: Preliminary analysis. (Left): Parameter size and MACs of modules. (Right): Sensitivity of pruning each module on LLaMA2-7B.
Coarse-grained Importance
Existing pruning work usually assigns the same sparsity to each block, but it is suboptimal for sparsity allocation. In fact, the blocks perform different functions and their importance varies significantly (Gromov et al., 2024b). Due to the residual structure, the effect of each block can be viewed as a transformation of the input representations.
Thus, we measure the coarse-grained importance of blocks $S_{g}$ through the saliency of transformation of feature activations during the forward process. Specifically, the $S_{g}$ of $l$-th block $\mathbf{B}^{\ell}$ is calculated as:
$S_{g}(\mathbf{B}^{\ell})$ $=\sum_{i=0}D({\bm{x}}^{\ell}_{i},{\bm{x}}^{\ell+1}_{i})$ (1)
$D({{\bm{x}}}^{\ell}_{i},{{\bm{x}}}^{\ell+1}_{i})$ $=\frac{1}{\pi}\arccos(\frac{{{\bm{x}}}^{\ell}_{i}\cdot{{\bm{x}}}^{\ell+1}_{i}}{\|{{\bm{x}}}^{\ell}_{i}\|\|{{\bm{x}}}^{\ell+1}_{i}\|})$ (2)
where ${\bm{x}}^{\ell}$ and ${\bm{x}}^{\ell+1}$ represent the input and output activation states of the $\ell$-th block. $D(\cdot)$ can be various distance measurements of two representations. Here we select the angular distance because it performs better than the others in our experiments. We then normalize $\mathbf{B}^{\ell}$ with the $\operatorname{Sigmoid}$ as:
$\operatorname{Norm}(S_{g}(\mathbf{B}^{\ell}))$ $=\operatorname{Sigmoid}(S_{g}(\mathbf{B}^{\ell})-\overline{S})$ (3)
$\operatorname{Sigmoid}(x)$ $=\frac{1}{1+e^{-\alpha\cdot x}}$ (4)
where $\overline{S}$ is the average importance score of all blocks. The function $\operatorname{Sigmoid}$ is introduced to process the scores non-linearly, which can make the distinction between blocks more significant, and $\alpha$ controls the intensity of significance. Finally, we assign the sparsity budgets across blocks based on the normalized importance scores as:
$\operatorname{Sparsity}(\mathbf{B}^{\ell})=\frac{\text{Norm}(S_{g}(\mathbf{B}^{\ell}))\cdot\gamma\cdot n}{\sum_{\ell=1}^{n}\operatorname{Norm}(S_{g}(\mathbf{B}^{\ell}))}$ (5)
where $n$ is the number of blocks and $\gamma$ represents the whole sparsity budget of the model.
Figure 3: The structural dependencies of FFN in LLaMA3. The blue part corresponds to the minimum unit of structured pruning. The red box represents the relative size of a matrix element in its row or column.
Dimension Adjustment
The Equation 5 may assign irregularly shaped weight matrices that do not satisfy the multiples of 64 or 128, thus destroying the parallelism of the tensor cores on the GPU. To this end, we introduce a simple adjustment during pruning to adjust the final dimensions of the pruned blocks to multiples of 128. For the $l$-th block $\mathbf{B}^{l}$, the final dimension $dim_{f}^{l}$ are computed as:
$dim_{f}^{l}$ $=(\left\lfloor\frac{dim_{o}\times\operatorname{Sparsity}(\mathbf{B}^{i})+64}{128}\right\rfloor)\times 128$ (6)
where $dim_{o}$ is the intermediate dimensions of original dense model. We present ablation results in Section 3.5 that demonstrate that the adjustment of dimensions significantly accelerates inference speed on GPUs. Notably, this enhancement is achieved with only a minimal increase in parameters and no detrimental impact on performance.
Fine-grained Importance
After assigning the proper sparsity to each block, we then identify the importance of pruning units (intermediate dimensions) within the block. Figure 3 shows the structural dependencies of three matrices used in FFN ($W_{u}$, $W_{g}$ and $W_{d}$): removing the intermediate dimensions is equivalent to subtracting the corresponding columns of $W_{u},W_{g}$ and the corresponding rows of $W_{d}$. The weight matrix represents the connections between neurons, where each row or column of weights influences the same neuron, implying that the fine-grained importance of weights is also related to their respective row or column. For the $i$-th intermediate dimension, we utilize the activation of $\mathbf{X}_{d}$ of $W_{d}$ and all weight matrices to calculate fine-grained importance score $S_{l}^{i}$:
$S_{l}^{i}$ $=F_{l}^{i}\cdot\left\|{\mathbf{X}}_{d}^{i}\right\|$ (7)
$F_{l}^{i}$ $=\sum_{j}(\frac{W^{ij}_{d}\cdot\left\|{\mathbf{X}}^{i}_{d}\right\|}{W^{*j}_{d}\cdot\left\|{\mathbf{X}}^{*}_{d}\right\|}+\frac{W^{ij}_{u}}{W^{i*}_{u}}+\frac{W^{ij}_{g}}{W^{i*}_{g}})$ (8)
[…]
2.3 Importance-guided Recovery Fine-tuning
In addition to the single-shot pruning scenario, we also explore the integration of recovery fine-tuning to further enhance performance at high sparsity. Our recovery setting follows Ma et al. (2023) to fine-tune with LoRA (Hu et al., 2022). Unlike the original LoRA, we propose an importance-guided method that adaptively assigns additional trainable parameters across different blocks. Specifically, for the $l$-th block, the rank $r^{l}$ of LoRA is determined based on the coarse-grained importance scores computed during pruning:
$r^{l}=\frac{\operatorname{Norm}(S_{g}(\mathbf{B}^{\ell}))\cdot\bar{r}\cdot n}{\sum_{\ell=1}^{n}\operatorname{Norm}(S_{g}(\mathbf{B}^{\ell}))}$ (9)
where $\bar{r}$ is the average rank allocated budget. In our experiments, we find that our recovery method is more efficient, achieving comparable performance while requiring less training data.
Recovery Data
We explore various datasets for recovery fine-tuning. We find that the quality and diversity of knowledge in the data are critical for recovery performance, especially on challenging tasks (e.g., MMLU). The details of datasets and results can be found in Appendices A.2 and B.2.
Sparsity Method WinoGrande PIQA HellaSwag OBQA ARC-e ARC-c MMLU FreebaseQA Average
0% LLaMA3-8B 72.93 80.96 79.17 45.00 77.90 53.16 62.09 72.62 67.98
20% w/o recovery
Magnitude-SP 50.99 51.31 26.18 30.20 25.43 25.76 23.71 0.52 29.26
Wanda-SP 67.56 75.41 65.99 42.00 65.40 41.38 46.20 39.11 55.38
FLAP 65.67 74.65 62.41 40.20 61.36 35.15 41.39 34.58 51.93
CFSP (ours) 70.32 77.64 72.74 41.20 68.10 43.86 56.43 38.59 58.61
50% w/o recovery
Magnitude-SP 51.14 50.27 26.47 29.40 25.08 26.49 23.08 0.52 29.06
Wanda-SP 59.51 63.98 45.71 33.00 44.95 27.99 27.08 6.15 38.54
FLAP 58.80 62.35 41.89 31.00 40.28 26.11 24.03 4.55 36.12
CFSP (ours) 62.04 66.76 49.96 31.80 48.74 30.89 32.39 10.83 41.67
w/ recovery
Wanda-SP 61.48 70.89 60.20 37.60 60.86 36.43 35.54 11.89 46.86
CFSP (ours) 65.51 72.03 61.45 36.20 62.37 37.54 40.37 18.32 49.22
Table 1: Zero-shot performance of pruned models on LLaMA3-8B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold indicates the best results under the same setting.
Sparsity Method WinoGrande PIQA HellaSwag OBQA ARC-e ARC-c MMLU FreebaseQA Average
0% LLaMA3-70B 80.35 84.71 48.60 84.93 85.90 64.16 75.36 81.53 75.69
20% w/o recovery
Magnitude-SP 51.93 58.38 32.69 29.40 32.41 27.73 24.44 0.52 32.19
Wanda-SP 77.19 82.92 82.50 49.20 81.65 58.28 66.74 79.65 72.27
FLAP 77.51 82.48 80.41 47.40 78.49 55.12 65.88 79.02 70.79
CFSP (ours) 80.66 83.51 83.97 46.40 83.46 61.43 73.04 80.18 74.08
50% w/o recovery
Magnitude-SP 51.22 52.72 27.04 30.00 25.46 26.62 23.52 0.55 29.64
Wanda-SP 73.95 76.44 73.80 44.00 66.79 43.94 54.91 42.26 59.51
FLAP 72.85 76.82 68.05 42.80 66.54 45.05 53.90 38.41 58.05
CFSP (ours) 75.06 78.89 75.95 43.60 71.34 46.67 59.74 46.42 62.20
w/ recovery
Wanda-SP 76.33 80.02 79.73 47.20 73.10 47.26 59.98 48.20 63.98
CFSP (ours) 78.30 81.01 80.18 45.20 76.65 51.54 65.52 54.77 66.65
Table 2: Zero-shot performance of pruned models on LLaMA3-70B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold indicates the best results under the same setting.
Sparsity Method WinoGrande PIQA OBQA HellaSwag ARC-e ARC-c MMLU FreebaseQA Average
0% LLaMA2-13B 72.22 80.52 45.20 79.38 77.48 49.06 50.51 67.57 65.24
20% w/o recovery
Magnitude-SP 49.96 60.01 25.60 39.89 42.93 29.86 25.51 0.65 34.30
Wanda-SP 70.01 78.45 43.00 73.87 72.56 44.28 41.70 40.69 58.07
FLAP 68.27 77.58 41.40 72.58 67.47 42.58 41.15 28.63 54.96
CFSP (ours) 71.75 78.29 43.60 75.76 73.48 47.27 46.99 54.65 61.47
[…]
Magnitude-SP 50.75 50.16 24.20 26.17 27.19 25.85 25.20 0.53 28.76
Wanda-SP 64.80 71.76 38.00 57.36 59.09 37.80 27.00 3.20 44.87
FLAP 60.54 68.50 36.60 53.95 48.78 30.97 23.00 1.50 40.48
CFSP (ours) 64.17 71.98 39.40 60.28 62.33 38.05 28.24 3.65 46.01
w/ recovery
Wanda-SP 66.85 74.37 40.60 67.15 68.31 41.04 35.24 25.35 52.36
CFSP (ours) 67.17 74.88 40.60 68.60 69.23 40.87 36.41 25.78 52.94
Table 3: Zero-shot performance of pruned models on LLaMA2-13B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold indicates the best results under the same setting.
3 Experiments
3.1 Experimental Setup
In this work, we target to prune intermediate dimensions of FFN in LLM and conduct experiments primarily on widely used LLM models: LLaMA3-{8B,70B} (Meta, 2024) and a middle-size LLaMA2-13B (Touvron et al., 2023b). We also conduct experiments on the latest LLaMA3.1-8B (Dubey et al., 2024) and more models from LLaMA family in Appendix B.1.
Evaluation Benchmarks
Following previous work (Sun et al., 2023; An et al., 2024), we evaluate the zero-shot performance of models across 5 well-known tasks: WinoGrande (Sakaguchi et al., 2020), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018). Since Jaiswal et al. (2023a) shows that LLM pruning methods tend to have significantly degraded performance on knowledge-intensive tasks, we also include two challenging QA tasks for zero-shot evaluation: MMLU (Hendrycks et al., 2021) and FreebaseQA (Jiang et al., 2019), which focus on factual knowledge. For language modeling performance, we evaluate models on WikiText2 (Merity et al., 2017). Following previous work, we use the LM-Evaluation-Harness (Gao et al., 2023) and LLM-Kick (Jaiswal et al., 2023a) with default hyperparameters for the corresponding tasks. More details of the evaluation are shown in Appendix A.1
Implementation Details
For the pruning stage, the calibration data is randomly selected from the WikiText2 (Merity et al., 2017) training set. Unless otherwise stated, the calibration set consists of 128 samples and each has approximately 1024 tokens following Sun et al. (2023). In the recovery fine-tuning stage, the pruned models are trained on the 0.1B tokens from FineWeb-Edu (Lozhkov et al., 2024) dataset with the next-token prediction loss. We set the average rank budget of IG-LoRA at 8 following Ma et al. (2023). More details and ablation of the implementation are shown in Appendix A.2 and Appendix B.2, respectively.
Baselines
We compare the single-shot pruning performance [^2] Without recovery fine-tuning.of CFSP against the following baselines: Magnitude-SP measures the importance criterion based on the magnitude of weights (Han et al., 2016; Jaiswal et al., 2023b). This baseline employs uniform sparsity across blocks. Wanda-SP is extended by the unstructured pruning method Wanda (Sun et al., 2023), which modifies the target pruning units to structured weights. We globally sort the pruning units across all blocks to identify redundant components, as this strategy tends to achieve better performance compared to adopting a local manner for individual blocks. FLAP (An et al., 2024) uses the stability of activations as an importance criterion, also applying a global sorting strategy. Notably, for a fair comparison, all baselines are implemented to prune intermediate dimensions of the FFN, which are the same as CFSP. Details are shown in Appendix A.3.
3.2 Main Results
Zero-shot Tasks
We present a performance comparison of the LLaMA3 family in Tables 1 and 2, as well as LLaMA2-13B in Table  3. In the single-shot pruning setting, CFSP consistently demonstrates superior average performance compared to baselines across various models at both 20% and 50% sparsity. Remarkably, CFSP achieves a promising accuracy of 32.39 on MMLU with 50% sparsity on LLaMA3-8B, while other baselines regress to chance-level accuracy (~25.0). This result underscores the potential of CFSP to perform well on more challenging tasks without retraining, even at high sparsity. Furthermore, CFSP is more favorable for larger models. At the 20% and 50% sparsity on LLaMA3-70B, CFSP maintains 97.9% and 82.2% of the original performance on average, respectively. We further experiment CFSP with recovery fine-tuning at 50% sparsity for each model. For comparison, we choose Wanda-SP, as it has the second-best average performance in single-shot pruning. We fine-tune pruned models with our proposed IG-LoRA on 0.1B tokens from the FineWeb-Edu dataset. We find that after recovery training, both pruning models are improved, especially on complex knowledge-sensitive tasks. CFSP still outperforms Wanda-SP in general, indicating the effectiveness of our proposed pruning and recovery approach.
Sparsity Method LLaMA3-8B LLaMA3-70B
0% Dense 6.82 5.26
w/o recovery
20% Wanda-SP 9.39 7.86
FLAP 9.40 8.21
CFSP(ours) 8.97 8.02
w/o recovery
50% Wanda-SP $19.49$ $13.53$
FLAP $21.06$ $13.37$
CFSP(ours) 17.45 13.02
w/ recovery
Wanda-SP $14.52$ $11.75$
CFSP(ours) 12.55 10.92
Table 4: Perplexity of pruning methods for LLaMA3-8B and LLaMA3-70B on WikiText2 validation set.
Language Modeling
Table 4 presents the perplexity on WikiText2. CFSP consistently achieves better results than baselines, except for the 20% sparsity on LLaMA3-8B, where it performs slightly worse than Wanda-SP. Additionally, the benefits of CFSP are more pronounced at higher sparsity.
Model Parameters Memory MACs Speed-up
CPU GPU
LLaMA3-8B 8.03B 16.06GB 3.64T 1.0x 1.0x
+ CFSP 5.21B 10.42GB 2.19T 2.3x 1.6x
LLaMA2-7B 6.73B 12.61GB 3.38T 1.0x 1.0x
+ CFSP 4.57B 8.62GB 2.17T 2.1x 1.5x
Table 5: Comparison of parameter size, memory usage, MACs, and inference speed-up on CPU/GPU. The pruned models (+CFSP) are under 50% sparsity.
3.3 Efficiency Evaluation
We assess the inference efficiency of the pruned models. The details of evaluation are shown in Appendix A.1. The results of 50% sparsity are shown in Table 5. Compared to the original dense models, CFSP reduces the parameters, memory, and MACs by 40% and achieves a speed-up over 1.5×1.5$\times$1.5 × on CPU and GPU. We also report the pruning and recovery time in Appendix A.2. In general, CFSP significantly improves efficiency, indicating its effectiveness for practical deployments of LLM.
Setting PPL↓ HellaSwag MMLU
coarse-grained importance ablation
(a) Uniform 9.08 70.84 50.91
(b) Euclidean 9.11 70.11 50.19
(c) Cosine 8.98 72.52 55.79
Angular (Ours) 8.97 72.74 56.43
fine-grained importance ablation
(d) Wanda 9.03 71.93 55.33
Eq (8) (Ours) 8.97 72.74 56.43
Table 6: Ablation of importance criterion of CFSP on LLaMA3-8B under 20% sparsity.
3.4 Ablation Study
Importance Criterion
We explore the effects of each component incorporated in the importance criterion of the proposed CFSP. Table 6 shows the ablation results under 20% sparsity of LLaMA3-8B. We first investigate the coarse-grained importance of blocks by comparing variants including: (a) uniform sparsity for each block, (b) Euclidean distance, or (c) cosine similarity as the coarse-grained importance criterion to allocation sparsity budget across blocks. As illustrated in Table 6, applying uniform sparsity or using Euclidean distance results in a notable performance decrease, particularly for zero-shot tasks. The angular distance (Eq. 2) used in CFSP achieves the best performance across tasks. For fine-grained importance ablation, as shown in the Table 6, the criterion outlined in Eq. 8 also demonstrates superior performance compared to the criterion utilized in Wanda.
Recovery Fine-tuning
To assess the impact of our proposed IG-LoRA for recovery, we compare it with the original LoRA. Figure 4 shows that IG-LoRA exhibits better performance than LoRA across various recovery data sizes and rank configurations. Furthermore, IG-LoRA achieves a performance comparable to LoRA trained on the full dataset while utilizing only 60% of data, highlighting the efficiency of IG-LoRA.
Figure 4: Results of different recovery fine-tuning methods at different data sizes.$r=$ 8/32 means the average rank budget configuration is set to 8 or 32.
Figure 5: Performance comparison of MMLU task between CFSP and Wanda-SP on LLaMA3-8B with various sparsity.
3.5 Analysis
Performance with Various Sparsity
Figure 5 presents the MMLU results of pruned models with sparsity from 5% to 50%. Under lower sparsity (10%), Wanda-SP is comparable to CFSP. As the sparsity increases, its performance decreases significantly, while CFSP still maintains promising performance even at 50% sparsity.
Figure 6: The effect of dimension adjustment. The speed-up is evaluated on a NVIDIA A800-80G.
Impact of Dimension Adjustment
Figure 6 compares the inference speed-up of whether to perform dimension adjustment during pruning. We observe that adjusting the intermediate dimension significantly accelerates models (1.6×1.6$\times$1.6 ×). However, without the adjustment, the latency of pruned models is comparable to the original dense models. Furthermore, the cost of adjustment is negligible and does not impact performance. For instance, on LLaMA3-8B, the number of parameters increased only by 0.43% after adjustment (from 5.21B to 5.23B). The average zero-shot performance remains comparable to that without adjustment (41.67 vs 41.61).
Visualization
Figure 7: The visualizations of normalized coarse-grained importance scores of each block on LLaMA3-8B and LLaMA3-70B.
We present a visualization of block importance scores on LLaMA3 in Figure 7. We find that the scores vary significantly across blocks: the first and last blocks exhibit the highest scores, whereas the intermediate blocks show lower scores. The varying importance explains why the uniform sparsity allocation is suboptimal and indicates that intermediate blocks exhibit greater redundancy, allowing for more aggressive pruning, while other blocks need to retain more weights.
4 Related Work
LLM Compression
The enormous computations of LLMs has prompted efforts in improving their efficiency, including quantization (Dettmers et al., 2022; Yao et al., 2022; Lin et al., 2023; Frantar et al., 2022; Xiao et al., 2023; Dettmers et al., 2023a, b; Shao et al., 2023; Xu et al., 2024b), distillation (Wang et al., 2022; Jiang et al., 2023; Wang et al., 2023; Hsieh et al., 2023; Agarwal et al., 2023; Gu et al., 2023) and KV cache compression (Sheng et al., 2023; Ge et al., 2023; Zhang et al., 2023b; Liu et al., 2024; Hooper et al., 2024; DeepSeek-AI et al., 2024). Pruning is another crucial method by eliminating redundant parameters. Most of the previous pruning work follows the unstructured pruning, which removes individual parameters according to their importance (Frantar and Alistarh, 2023; Sun et al., 2023; Dettmers et al., 2023b; Xu et al., 2024a; Zhang et al., 2024b). However, this paradigm requires specialized hardware support to speed up. In contrast, structured pruning eliminates the structural group of weights, facilitating a more convenient deployment on general hardware (Wang et al., 2020; Xia et al., 2022). Some work proposes to remove redundant layers in LLMs (Men et al., 2024; Yang et al., 2024b; Gromov et al., 2024b), while dropping entire layers leads to a significant performance drop. For pruning on more fine-grained units, some work formulates pruning as a constrained optimization problem by introducing learnable masks to search (Xia et al., 2023; Dery et al., 2024; Muralidharan et al., 2024; Gunter et al., 2024). Zhang et al. (2024a) performs iteratively to prune the coupled weights until the desired sparsity is achieved. Ma et al. (2023) and Zhang et al. (2023a) use gradient information to guide pruning. These methods incur substantial pruning overhead, particularly in the case of large-scale models. An et al. (2024) eliminates channels based on their activation fluctuations using only forward passes. In this work, we also aim to achieve efficient structured pruning using only forward passes.
Sparsity in Transformer
Sparsity is a common trait in neural networks (Allen-Zhu et al., 2019; Frankle and Carbin, 2019; Jaszczur et al., 2021) and a lot of work explores sparsity in Transformer, such as attention (Voita et al., 2019; Michel et al., 2019; Hao et al., 2021; Zhu et al., 2021) or FFN (Wang et al., 2020; Zhang et al., 2022; Zuo et al., 2022). The dynamic sparsity has also garnered attention (Liu et al., 2023; Wang et al., 2024), which adaptively selects a portion of the model based on input. Yin et al. (2024) find that non-uniform sparsity yields better results for LLM unstructured pruning, which is consistent with our observation in structured pruning.
5 Conclusion
In this work, we explore structured pruning for Large Language Models (LLMs). We propose an efficient pruning framework named CFSP, which leverages coarse to fine-grained activation information as an importance criterion to determine the redundant parts to prune. For the coarse-grained importance, we measure the saliency of transformations of each block and use this criterion to assign the sparsity budget across blocks. For weights within each block, we utilize a fine-grained criterion to remove redundant parts to obtain compact models. We also introduce an efficient recovery fine-tuning method IG-LoRA that adaptively assigns additional trainable parameters based on the importance of blocks. Extensive experimental results demonstrate that CFSP outperforms existing methods across various models and sparsity levels, both in single-shot pruning and in recovery fine-tuning. Meanwhile, even at high sparsity, our method can maintain promising performance on challenging tasks such as MMLU and FreebaseQA compared to the original dense models.
Limitations
CFSP is a fast and efficient structured pruning method for large language models (LLMs), while it also has some limitations. First, our experiments focus on the LLaMA family of models (Touvron et al., 2023a, b; Meta, 2024; Dubey et al., 2024), as they are among the most advanced open-source LLMs currently. We will extend our method to a broader range of models in the future. Additionally, we do not prune attention heads, as this has been shown to cause significant performance degradation, especially for models that have grouped query attention (GQA) (Ainslie et al., 2023) like LLaMA3. Further research is needed to develop more effective pruning strategies, especially in the context of attention optimization techniques like GQA.
Appendix A Details of Experimental Setup
A.1 Details of Evaluation Benchmarks
Zero-shot Tasks Evaluation
In this work, we consider the following tasks for evaluating zero-shot performance, along with their respective evaluation metrics: WinoGrande (Sakaguchi et al., 2020) with the accuracy, PIQA (Bisk et al., 2020) with the normalized accuracy, OBQA (Mihaylov et al., 2018) with the normalized accuracy, HellaSwag (Zellers et al., 2019) with the normalized accuracy, ARC-easy/challenge (Clark et al., 2018) with the normalized accuracy, MMLU (Hendrycks et al., 2021) with the accuracy, FreebaseQA (Jiang et al., 2019) with the exact-match score. The first 6 tasks are general common sense reasoning tasks, while the others are knowledge-intensive. We evaluate WinoGrande, PIQA, HellaSwag, BoolQ, ARC-e/c, and MMLU by LM-Evaluation-Harness (Gao et al., 2023) in multiple choice form: we compute the loglikelihood for each choice and report the accuracy for the highest choice. For FreebaseQA, the evaluation is run with LLM-Kick (Jaiswal et al., 2023a).
Language Modeling Evaluation
We evaluate language modeling performance on WikiText2 (Merity et al., 2017) validation set with the setting of (Gao et al., 2023). The input length is 1024 for the 8B/13B models and 256 for the 70B models.
Inference Efficiency Evaluation
We evaluate the speed-up of CPU on an Intel Xeon E5-466 2640 v4 CPU and the speed-up of GPU on a single A800-80G GPU. We set the sequence length to 1024 and the batch size to 1.
A.2 Implementation Details
We implement CFSP with Huggingface Transformer (Wolf et al., 2019). We perform experiments on NVIDIA A800-80G GPUs. The pruning stage is conducted on 1 GPU for the 7B/13B models and 8 GPUs for the 70B models. Unless otherwise stated, the calibration dataset consists of 128 samples and each has approximately 1024 tokens following Sun et al. (2023); An et al. (2024). By default, for the 7B/8B/13B models, $\alpha$ in Equation 3 is set to 1, whereas for the 70B model, it is set to 3. For the recovery fine-tuning stage, the average rank budget is set to 8 by default. We explore the following datasets for recovery training:
In our experiments, we use FinWeb-Edu as our default recovery data since it achieves the best performance across all tasks. More experimental results are shown in Appendix B.2. We use the AdamW optimizer with a learning rate of 2e-4 for the 8B and 13B models, and 1e-4 for the 70B models. The batch size is set to 128. We use 8 GPUs to fine-tune the pruned 7B/8B/13B models and 16 GPUs for the 70G models. We also show the details of time cost for pruning and recovery in Table 7.
Model Size prune recovery
device time device time
7/8B 1xA800-80G 2min 8xA800-80G 0.5h
13B 1xA800-80G 4-5min 8xA800-80G 0.92h
70B 8xA800-80G 15min 16xA800-80G 5.42h
Table 7: Details of time cost of pruning and recovery for different model sizes.
A.3 Details of Baselines
In this section, we present more details of baselines in comparison:
Magnitude-SP
measures the importance criterion based on the magnitude of weights (Han et al., 2016; Jaiswal et al., 2023b). This baseline employs with uniform sparsity across blocks.
Wanda-SP
is extended by the unstructured pruning method Wanda (Sun et al., 2023), which modifies the target pruning units to structured weights. This baseline uses the product of weights and activations as an importance criterion. We globally sort the pruning units across all blocks to identify redundant components, as this strategy tends to achieve better performance compared to adopting a local manner for individual blocks.
FLAP
(An et al., 2024) is a training-free structured pruning method for LLM, using the stability of activations as an importance criterion with a global sorting strategy. We follow its optimal setting: Weighted Input Feature Variance.
For a fair comparison, all baselines are implemented to prune the intermediate dimensions of FFN, which are the same as CFSP. Since the original FLAP paper only reports the results of both MHA and FFN pruning on LLaMA, we reimplement based on their official code and conduct on more models.
Sparsity Method WinoGrande PIQA OBQA HellaSwag ARC-e ARC-c MMLU Average
0% LLaMA3.1-8B 73.95 81.01 44.8 78.91 80.85 53.33 62.95 69.74
20% w/o recovery
Wanda-SP 67.96 74.65 41.00 65.95 64.44 39.85 45.21 57.01
FLAP 64.64 73.72 40.60 61.92 61.11 35.92 38.20 53.73
CFSP (ours) 71.51 76.88 41.60 72.28 70.39 44.88 54.59 63.59
50% w/o recovery
Wanda-SP 58.88 63.76 32.20 46.03 46.93 29.52 26.39 43.39
FLAP 58.09 59.96 31.00 41.98 39.56 26.02 23.15 39.97
CFSP (ours) 61.09 66.16 32.40 49.31 48.70 29.95 32.05 45.67
w/ recovery
Wanda-SP 61.88 70.78 36.80 59.58 61.53 36.43 36.44 51.92
CFSP 65.19 71.16 36.40 61.23 62.54 37.29 40.65 55.83
Table 8: Zero-shot performance of pruned models on LLaMA3.1-8B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold results indicate the best results under the same setting.
Sparsity Method WinoGrande PIQA OBQA HellaSwag ARC-e ARC-c MMLU FreebaseQA Average
0% LLaMA2-7B 69.06 79.11 44.20 76.02 74.62 46.33 41.25 68.39 62.37
20% w/o recovery
Magnitude-SP 48.70 52.12 24.40 28.77 30.18 24.32 25.84 0.55 29.36
Wanda-SP 66.93 76.50 41.80 70.82 64.48 38.57 32.19 44.44 54.46
FLAP 65.51 75.84 40.00 69.64 60.82 37.29 30.86 28.65 51.07
CFSP (ours) 67.25 76.88 40.60 72.05 68.77 41.47 36.33 38.99 55.29
[…]
Magnitude-SP 50.20 48.20 27.00 26.32 26.52 29.10 26.84 0.53 29.34
Wanda-SP 61.56 66.49 34.80 52.23 45.37 28.07 25.45 4.15 39.76
FLAP 57.62 66.00 32.80 48.09 40.07 27.39 23.04 0.90 36.99
CFSP (ours) 61.64 67.36 35.20 53.96 48.61 30.20 23.07 4.35 40.55
w/ recovery
Wanda-SP 63.85 71.11 37.80 61.40 57.08 35.04 26.32 20.90 46.69
CFSP (ours) 65.11 70.73 37.00 61.96 58.88 36.26 29.46 20.05 47.43
Table 9: Zero-shot performance of pruned models on LLaMA2-7B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold indicates the best results under the same setting.
Sparsity Method WinoGrande PIQA OBQA HellaSwag ARC-e ARC-c MMLU Average
0% LLaMA-7B 70.09 79.16 44.06 76.21 72.85 44.80 29.92 59.58
20% w/o recovery
Magnitude-SP 49.33 52.12 24.20 27.20 28.66 25.68 24.85 33.15
Wanda-SP 67.88 76.17 41.00 70.54 66.67 39.85 27.63 55.68
FLAP 66.61 75.63 42.00 68.91 66.33 38.82 26.95 55.04
CFSP (ours) 68.43 75.90 41.20 71.48 69.44 42.66 27.75 56.69
50% w/o recovery
Magnitude-SP 52.01 49.24 26.20 26.31 26.43 26.96 24.87 33.15
Wanda-SP 63.30 65.38 37.00 52.13 47.81 29.10 24.16 45.55
FLAP 60.14 65.56 36.00 50.23 44.82 29.01 24.46 44.32
CFSP (ours) 63.69 66.21 37.20 54.55 47.98 30.12 24.03 46.25
w/ recovery
Wanda-SP 65.51 71.33 38.20 61.29 58.42 34.04 24.53 50.47
CFSP 65.55 71.22 39.20 61.31 58.96 34.73 25.35 50.90
Table 10: Zero-shot performance of pruned models on LLaMA-7B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold indicates the best results under the same setting.
Sparsity Method WinoGrande PIQA OBQA HellaSwag ARC-e ARC-c MMLU Average
0% LLaMA-13B 72.69 80.20 44.80 79.08 74.71 47.70 41.24 62.92
20% w/o recovery
Magnitude-SP 48.86 58.43 27.40 33.29 33.80 29.10 23.36 36.32
Wanda-SP 70.56 77.53 41.40 75.40 66.25 41.98 31.65 57.82
FLAP 70.56 77.09 41.40 74.19 68.77 43.60 31.77 58.19
CFSP (ours) 71.43 78.13 42.80 76.31 68.81 45.14 38.07 60.10
[…]
Magnitude-SP 50.99 50.98 26.20 27.16 27.27 27.30 23.94 33.40
Wanda-SP 67.01 67.46 37.00 61.44 49.83 31.14 27.23 48.73
FLAP 64.24 70.24 36.00 56.73 52.86 33.19 25.23 48.35
CFSP (ours) 68.43 71.76 37.60 63.50 59.85 37.46 27.38 52.28
w/ recovery
Wanda-SP 67.48 75.08 39.20 68.14 39.59 64.48 31.97 55.13
CFSP (ours) 68.82 74.76 41.40 68.85 41.30 67.13 35.63 56.84
Table 11: Zero-shot performance of pruned models on LLaMA-13B under 20% and 50% sparsity. For 50% sparsity, we also show the results after recovery fine-tuning. Bold indicates the best results under the same setting.
Appendix B More Results and Analysis
B.1 More Results on LLaMA
Results of LLaMA3.1-8B
In addition to the experiments presented in Section 3.2, we also conduct experiments on the latest powerful model LLaMA3.1-8B. The results are shown in Table 8. In the single-shot pruning setting, CFSP consistently outperforms other baselines across a variety of tasks and sparsity budgets. Furthermore, we experiment with recovery fine-tuning for CFSP and Wanda-SP. As observed with previous models, our method still achieves better results.
Results of LLaMA2-7B
Table 9 shows the results on LLaMA2-7B. In the single-shot pruning setting, CFSP consistently exceeds other baselines on various tasks and sparsity levels. Additionally, we perform recovery fine-tuning for both CFSP and Wanda-SP. As with other models, CFSP also provides superior performance.
Results of LLaMA1
Since the LLaMA1 family models were released earlier and no longer the best open-source LLMs, we do not include their results in Section 3.2. Here, we present the zero-shot performance comparison of LLaMA1 family in Table 10 and Table 11777 The results of Wanda-SP reported by us differ from those in An et al. (2024) since we employ a global sorting strategy as described in Appendix A.3.. It can be observed that on LLaMA-7B and LLaMA-13B, CFSP consistently achieves the best average performance at different sparsity. An interesting phenomenon is that on some challenging tasks (e.g. MMLU), all pruning methods exhibit performance close to chance-level accuracy at 50% sparsity. Compared to the results on the LLaMA3 herd of models, this could be attributed to the LLaMA1 family models’ inherently weaker performance on these tasks, with high-sparsity pruning further degrading this aspect of their performance.
Figure 8: The impact of the size of calibration data. The models are pruned from LLaMA3-8B under 50% sparsity.
B.2 More Analysis of CFSP
Figure 9: The effect of hyperparameter $\mathbf{\alpha}$ in calculating block importance. The models are pruned under 50% sparsity.
Model Datasets WinoGrande PIQA OBQA HellaSwag ARC-e ARC-c MMLU Average
LLaMA2-7B Slimpajama 64.01 70.40 36.40 60.65 56.78 33.36 24.95 47.46
Alpaca-cleaned 64.33 70.24 37.60 61.22 59.30 34.47 24.33 50.21
Knowledge-pile 64.80 70.13 36.80 60.45 58.42 34.81 24.55 49.99
FineWeb-edu 65.11 70.73 37.00 61.96 58.88 36.26 29.46 51.34
LLaMA3-8B Slimpajama 64.17 70.95 35.00 60.62 57.03 33.62 37.55 51.28
Alpaca-cleaned 59.67 67.85 34.80 60.97 57.41 35.24 37.89 50.55
Knowledge-pile 66.14 71.38 35.60 60.89 62.05 36.69 38.07 52.97
FineWeb-edu 65.51 72.03 36.20 61.45 62.37 37.54 40.37 53.64
Table 12: Zero-shot performance of various datasets for recovery fine-tuning. All methods are trained with the same tokens (0.1B). Bold indicates the best results on each model.
Impact of Hyperparameter $\mathbf{\alpha}$
In Equation 3, we introduce a hyperparameter $\alpha$ to control the intensity of significance during calculating block importance. In preliminary experiments, we explore the impact of different $\alpha$ and the results are shown in Figure 9. We observe that for smaller models like LLaMA3-8B, a smaller $\alpha$ is better, while for larger model like LLaMA3-70B model, a larger $\mathbf{\alpha}$ is more appropriate. Finally, we set $\alpha=1$ for the 7B/8B/13B models and $\alpha=3$ for the 70B models.
Impact of Calibration Data Sizes
We investigate the impact of calibration data sizes. Figure 8 presents the results of 3 tasks on LLaMA3-8B with 20% sparsity. We find that the data with 128 examples yield the best overall performance.
Impact of Recovery Data
As described in Appendix A.2, we explore various datasets for recovery fine-tuning. As shown in Table 12, FineWeb-Edu consistently outperforms others in a variety of tasks, particularly demonstrating significant improvements in knowledge-intensive tasks such as MMLU and FreebaseQA, which is shown challenging for pruned models (Jaiswal et al., 2023a). Thus, we select it for recovery fine-tuning.